### v3

# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

Abstract
Integrating structured graph data with rich textual information from nodes poses a significant challenge, particularly for heterophilic node classification. Current approaches often struggle with computational costs or effective fusion of disparate modalities. We propose Graph Masked Language Model (GMLM), a novel architecture efficiently combining Graph Neural Networks (GNNs) with Pre-trained Language Models (PLMs). GMLM introduces three key innovations: (i) a dynamic active node selection strategy for scalable PLM text processing; (ii) a GNN-specific contrastive pretraining stage using soft masking with a learnable graph [MASK] token for robust structural representations; and (iii) a dedicated fusion module integrating RGCN-based GNN embeddings with PLM (GTE-Small & DistilBERT) embeddings. Extensive experiments on heterophilic benchmarks (Cornell, Wisconsin, Texas) demonstrate GMLM’s superiority. Notably, GMLM(DistilBERT) achieves significant performance gains, improving accuracy by over 4.7% on Cornell and over 2.0% on Texas compared to the previous best-performing baselines. This work underscores the benefits of targeted PLM engagement and modality-specific pretraining for improved, efficient learning on text-rich graphs.
1 Introduction
Graph Neural Networks (GNNs) [1, 2] have demonstrated remarkable success in node classification, link prediction, and graph-level tasks. However, many such architectures focus primarily on structural information and local neighborhood features, potentially overlooking rich semantic content embedded within node attributes, especially textual attributes.. While recent advances in natural language processing, particularly Pre-trained Language Models (PLMs) like BERT [3], have revolutionized our ability to capture contextual information from text, their effective and scalable integration with graph learning, especially for text-rich graphs, remains an active area of research. The challenge is particularly pronounced in heterophilic graphs, where connected nodes may exhibit dissimilar features and labels, requiring a nuanced understanding of both structure and semantics.
While architectures like Graph Attention Networks (GATs) [4] offer learnable attention mechanisms for neighborhood aggregation, and Relational Graph Convolutional Networks (RGCNs) [5] can handle typed edges, they may not inherently utilize deep semantic understanding from node-associated texts without explicit integration with PLMs. Concurrently, masked modeling techniques, central to the success of PLMs, have proven highly effective for learning contextual representations. This prompts the question: How can we adapt and combine the strengths of GNNs and PLM-style masked modeling to enhance representation learning on text-rich graphs, while addressing computational scalability ?
Traditional node feature masking in GNNs [6] often involves random feature perturbation or complete removal, which can disrupt learning or fail to provide sufficiently informative training signals. Furthermore, applying PLMs to all node texts in large graphs is often computationally prohibitive.
In this paper, we present Graph Masked Language Model (GMLM), a novel framework that bridges GNNs and PLMs for improved node representation learning on text-rich graphs, with a focus on heterophilic settings and computational efficiency. Our approach introduces three key innovations:
  1. (1)
    A dynamic active node selection strategy that stochastically identifies a subset of nodes during each training iteration. Only the texts associated with these active nodes are processed by the PLM, and the primary task loss is computed only on these nodes, significantly reducing computational overhead.
  2. (2)
    A GNN-specific contrastive pretraining stage that employs a soft masking mechanism. This mechanism interpolates original node features with a learnable graph-specific [MASK] token embedding, guided by a dynamically generated perturbation mask. This stage pre-trains the GNN (an RGCN in our implementation) to learn robust structural representations before end-to-end fine-tuning.
  3. (3)
    A dedicated fusion network designed to effectively combine the comprehensive structural embeddings derived from the GNN path with the selectively generated, deep semantic embeddings from the PLM path, leading to rich, multi-modal node representations.
This combination allows our model to learn both structural and semantic patterns effectively, leading to more robust and contextually aware node representations, particularly beneficial for challenging heterophilic node classification tasks.
Our main contributions in this work can be summarized as follows:
  1. 1.
    We propose a novel architecture, GMLM, that efficiently integrates a GNN (specifically an RGCN) with a PLM by introducing a dynamic active node selection mechanism for scalable text processing and a GNN-focused contrastive pretraining phase.
  2. 2.
    We introduce a GNN contrastive pretraining method utilizing a soft feature masking technique with a learnable graph [MASK] token. This pretrains the GNN to capture robust structural information and initializes the [MASK] token embedding meaningfully (Section 5.1).
  3. 3.
    We design a fusion strategy and an end-to-end fine-tuning process that utilizes the pre-trained GNN and selectively engages the PLM, allowing for effective learning from both graph structure and node texts for improved node classification, especially on heterophilic graphs (Section 5.2).
2 Related Works
Graph Neural Networks (GNNs) [2] provide a powerful framework for learning from graph-structured data. Building upon early spectral and spatial convolutional methods, Graph Attention Networks (GATs) [4] integrated self-attention [7]. This allows adaptive weighting of neighbors during message passing, enhancing model expressivity and suitability for irregular graph structures, as detailed in recent surveys.
Efforts to adapt the Transformer architecture, initially designed for sequences, to graph data aim to capture long-range dependencies via global self-attention, potentially overcoming limitations of traditional message passing [8], [9]. These Graph Transformers, reviewed systematically [10], can incorporate positional information, handle diverse graph scales, achieve competitive results, and utilize language model pretraining benefits on text-rich graphs.
Masked Pretraining Strategies, popularized by language models [3] that learn contextual representations via masked token prediction, have been adapted for graphs. The Masked Graph Autoencoder (MaskGAE) [11] framework, for instance, masks graph edges and trains the model to reconstruct them, thereby learning robust representations for downstream tasks like link prediction and node classification.
Integration of Language Models and Graph Structures seeks to combine textual and structural information. Key directions include: fusing pretrained language models (LMs) with GNNs for joint reasoning over text and structure [12]; injecting structural information directly into LMs using graph-guided attention mechanisms, potentially removing the need for separate GNN modules [13]; interpreting Transformers as GNNs operating on complete graphs, allowing explicit modeling of edge information [14]; and jointly training large LMs and GNNs to utilize both contextual understanding and structural learning capabilities [15]. We provide  Table 6 in  Appendix E giving an overview of the difference in methods and implementation.
3 Preliminaries
In this section, we introduce the fundamental concepts and notations used throughout this paper, including graph representation, Graph Neural Networks (specifically RGCNs), Pre-trained Language Models, and the concept of heterophily in graphs.
3.1 Graph Representation and Notation
A graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of $N=|\mathcal{V}|$ nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$ is the set of $M=|\mathcal{E}|$ edges. Each edge $(v_{i},r,v_{j})\in\mathcal{E}$ connects node $v_{i}$ to node $v_{j}$ under a relation type $r\in\mathcal{R}$, where $\mathcal{R}$ is the set of possible relation types. For graphs without explicit multiple relation types, we can consider a single relation type, effectively making it $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Each node $v_{i}\in\mathcal{V}$ is associated with an initial feature vector $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$, where $d_{x}$ is the input feature dimensionality. In the context of this work, nodes are also associated with textual descriptions $T=\{t_{1},t_{2},\dots,t_{N}\}$, where $t_{i}$ is the raw text for node $v_{i}$. The goal of node classification is to predict a label $y_{i}\in\mathcal{Y}$ for each node (or a subset of nodes), where $\mathcal{Y}$ is the set of possible labels.
3.2 Graph Neural Networks (GNNs)
Graph Neural Networks (GNNs) are a class of neural networks designed to operate directly on graph-structured data. They typically learn node representations by aggregating information from their local neighborhoods. A generic GNN layer can be expressed as:
$\mathbf{h}_{i}^{(l+1)}=\text{UPDATE}^{(l)}\left(\mathbf{h}_{i}^{(l)},\text{AGGREGATE}^{(l)}\left(\{(\mathbf{h}_{j}^{(l)},\mathbf{e}_{ij}):v_{j}\in\mathcal{N}(v_{i})\}\right)\right)$
where $\mathbf{h}_{i}^{(l)}$ is the representation of node $v_{i}$ at layer $l$, $\mathcal{N}(v_{i})$ is the set of neighbors of $v_{i}$, $\mathbf{e}_{ij}$ is the feature of the edge between $v_{i}$ and $v_{j}$ (if available), and AGGREGATE and UPDATE are differentiable functions ( sum, mean, max, or more complex neural networks).
3.2.1 Relational Graph Convolutional Networks (RGCNs)
Relational Graph Convolutional Networks (RGCNs) [5] extend GCNs to handle multi-relational graphs (heterogeneous graphs with different edge types). The message passing for a node $v_{i}$ in an RGCN layer is typically defined as:
$\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{v_{j}\in\mathcal{N}_{r}(v_{i})}\frac{1}{c_{i,r}}\mathbf{W}_{r}^{(l)}\mathbf{h}_{j}^{(l)}+\mathbf{W}_{0}^{(l)}\mathbf{h}_{i}^{(l)}\right)$ (2)
where $\mathcal{N}_{r}(v_{i})$ is the set of neighbors of node $v_{i}$ under relation $r\in\mathcal{R}$, $\mathbf{W}_{r}^{(l)}$ is a relation-specific learnable weight matrix for relation $r$ at layer $l$, $\mathbf{W}_{0}^{(l)}$ is a weight matrix for self-loops, $c_{i,r}$ is a normalization constant ( $|\mathcal{N}_{r}(v_{i})|$), and $\sigma$ is an activation function ( ReLU). To handle a large number of relations, RGCNs often use basis decomposition or block-diagonal decomposition for $\mathbf{W}_{r}^{(l)}$.
3.3 Pre-trained Language Models (PLMs)
Pre-trained Language Models (PLMs), such as BERT [3] and its variants ( DistilBERT [16]), have achieved state-of-the-art results on various natural language processing (NLP) tasks. These models are typically based on the Transformer architecture [17] and are pre-trained on massive text corpora using self-supervised objectives like Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). Given an input text sequence ( a node’s textual description $t_{i}$), a PLM first tokenizes it into a sequence of sub-word tokens $[w_{1},w_{2},\dots,w_{S}]$. These tokens are then converted into input embeddings (sum of token, position, and segment embeddings) and fed into multiple Transformer layers. The output of the PLM is a sequence of contextualized embeddings for each token. For sentence-level or document-level tasks, often the embedding of a special [CLS] token ( $\mathbf{h}_{[CLS]}$) or an aggregation ( mean pooling) of token embeddings is used as the representation of the entire input text. In our work, we utilize a PLM (specifically DistilBERT) to obtain semantic embeddings $\mathbf{h}_{PLM}(t_{i})$ from node texts $t_{i}$.
3.4 Node Feature Masking in GNNs
Masking is a common technique in self-supervised learning. In the context of GNNs, node feature masking aims to train the model to reconstruct or predict properties of masked nodes or their features, thereby learning robust representations. A simple approach might involve replacing a fraction of node features $\mathbf{x}_{i}$ with a zero vector or a special [MASK] token embedding before feeding them into the GNN [18, 19]. Our GNN contrastive pretraining phase builds upon this idea but introduces a soft masking mechanism with a learnable graph-specific [MASK] token.
3.5 Heterophily in Graphs
Graph homophily refers to the principle that nodes tend to connect to other nodes that are similar to themselves ( in terms of features or labels). Conversely, graph heterophily (or disassortativity) describes the scenario where connected nodes are often dissimilar [20]. Many real-world graphs exhibit heterophily, such as protein-protein interaction networks or certain social networks. Standard GNNs, which assume homophily by smoothing features over neighborhoods, may perform poorly on heterophilic graphs. Designing GNN architectures that can effectively handle heterophily is an ongoing research challenge, often requiring models to capture more complex relational patterns beyond simple neighborhood similarity. Our work aims to contribute to this by integrating rich textual semantics which can provide crucial context in such settings.
3.6 Contrastive Learning
Contrastive learning is a self-supervised learning paradigm that aims to learn representations by maximizing agreement between differently augmented "views" of the same data sample (positive pairs) while minimizing agreement between views of different samples (negative pairs). A common contrastive loss function is the NT-Xent loss [21]:
$\mathcal{L}_{NT\text{-}Xent}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1}^{2B}\mathbb{I}_{[k\neq i]}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
where $(\mathbf{z}_{i},\mathbf{z}_{j})$ is a positive pair of augmented representations from the same original sample within a batch of size $B$, $\text{sim}(\cdot,\cdot)$ is a similarity function ( cosine similarity), $\tau$ is a temperature hyperparameter, and the sum in the denominator is over all $2B$ augmented samples in the batch, excluding $\mathbf{z}_{i}$ itself. Our GNN pretraining phase utilizes this principle.
Figure 1: Overview of the GMLM architecture and its two-phase training procedure. (Phase A) GNN Contrastive Pre-Training: Two augmented views of the input graph features are generated using random masks and a learnable GNN [MASK] token, then processed by a shared RGCN stack. An NT-Xnet loss encourages similar representations for positive pairs. (Phase B) End-to-End Fine-Tuning: The graph and associated texts are input. Active Node Selection directs text processing for the PLM branch and guides GNN input feature soft masking. Outputs from the GNN and PLM branches are projected, fused, and passed to a classification head. The GNN is initialized from Phase A.
4 Methodology
We introduce GMLM (Graph Masked Language Model), a novel architecture designed to effectively integrate graph structural information with rich textual semantics from nodes for robust node representation learning, particularly targeting heterophilic graphs. The GMLM architecture is characterized by two main processing pathways for graph and text data, a specialized GNN input perturbation mechanism, and a dedicated fusion module. An overview of the entire pipeline is depicted in Figure 1.
4.1 GNN Input Feature Perturbation with Soft Masking
To enhance the robustness of the GNN and to provide a meaningful training signal during both the contrastive pretraining phase (detailed in Section 5.1) and the end-to-end fine-tuning phase, we employ a soft masking strategy for the input node features. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{x}}$ be the matrix of initial node features. Given a binary perturbation mask $\mathbf{m}_{pert}\in\{0,1\}^{N}$, which indicates the nodes whose features are to be perturbed, the soft masking operation produces perturbed features $\mathbf{X}^{\prime}$ as follows:
$\mathbf{x}^{\prime}_{i}=\begin{cases}(1-\beta)\cdot\mathbf{x}_{i}+\beta\cdot\mathbf{e}_{mask}^{gnn}&\text{if }m_{pert,i}=1\\ \mathbf{x}_{i}&\text{if }m_{pert,i}=0\end{cases}$ (4)
where $\mathbf{x}_{i}$ is the original feature vector for node $v_{i}$, $\mathbf{x}^{\prime}_{i}$ is the perturbed feature vector, $\mathbf{e}_{mask}^{gnn}\in\mathbb{R}^{d_{x}}$ is a learnable GNN-specific mask token embedding, and $\beta\in[0,1]$ is a hyperparameter controlling the interpolation weight. This GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$ is initialized using Xavier uniform initialization and learned alongside other model parameters. The perturbation mask $\mathbf{m}_{pert}$ is generated dynamically during training (details in Sections 5.1 and 5.2).
4.2 Graph Neural Network (GNN) Branch
[…]
To prepare for fusion with PLM embeddings, these raw GNN embeddings are projected to match the PLM’s hidden dimensionality $d_{PLM}$ using a linear projection layer:
$\mathbf{h}_{G,i}^{proj}=\mathbf{W}_{G}^{proj}\mathbf{h}_{G,i}+\mathbf{b}_{G}^{proj}$ (5)
where $\mathbf{W}_{G}^{proj}\in\mathbb{R}^{d_{PLM}\times d_{G}}$ and $\mathbf{b}_{G}^{proj}\in\mathbb{R}^{d_{PLM}}$ are learnable parameters. The activation for these projected embeddings occurs as part of the subsequent fusion network.
4.3 Pre-trained Language Model (PLM) Branch with Active Node Selection
The PLM branch processes the textual descriptions associated with nodes to capture rich semantic information. To manage computational costs, particularly for large graphs, we incorporate an Active Node Selection mechanism. During each training iteration (in the fine-tuning stage), a subset of nodes $\mathcal{V}_{active}\subseteq\mathcal{V}$ is dynamically selected based on an active node mask $\mathbf{m}_{active}\in\{0,1\}^{N}$ (generation details in Section 5.2). Only the texts $\{t_{i}:v_{i}\in\mathcal{V}_{active}\}$ are processed by the PLM.
For each selected node $v_{i}\in\mathcal{V}_{active}$, its text $t_{i}$ is tokenized and fed into a pre-trained language models  [^1] We choose these models as they are computationally efficient but also maintain almost equivalent performance to their larger variants.: GTE-small[22] and DistilBERT [16]. The PLM outputs contextualized embeddings, from which we extract a fixed-size representation, specifically the embedding of the [CLS] token:
$\mathbf{h}_{PLM,i}=\text{PLM-Encoder}(t_{i})_{[CLS]}\in\mathbb{R}^{d_{PLM}}\quad\forall v_{i}\in\mathcal{V}_{active}$
For nodes $v_{j}\notin\mathcal{V}_{active}$, their PLM embeddings are effectively zero vectors when constructing the full PLM embedding matrix $\mathbf{H}_{PLM}^{full}\in\mathbb{R}^{N\times d_{PLM}}$ for fusion.
4.4 Fusion Network and Classification Head
To integrate the structural information from the GNN branch and the semantic information from the PLM branch, we employ a fusion network. For each node $v_{i}$, its projected GNN embedding $\mathbf{h}_{G,i}^{proj}$ (Equation 5) is concatenated with its corresponding (potentially zero) PLM embedding $\mathbf{h}_{PLM,i}^{full}$ from $\mathbf{H}_{PLM}^{full}$:
$\mathbf{h}_{concat,i}=[\mathbf{h}_{G,i}^{proj};\mathbf{h}_{PLM,i}^{full}]\in\mathbb{R}^{d_{PLM}+d_{PLM}}$ (7)
This concatenated vector is then passed through a fusion network, which is a multi-layer perceptron (MLP) designed to learn a joint representation:
$\mathbf{h}_{fused,i}=\text{MLP}_{fusion}(\mathbf{h}_{concat,i})\in\mathbb{R}^{d_{fused}}$ (8)
In our model, $\text{MLP}_{fusion}$ consists of two linear layers interspersed with Layer Normalization, ReLU activation, and Dropout. It projects the $2\cdot d_{PLM}$ dimensional concatenated vector to an intermediate hidden dimension, and then to a $d_{fused}$ dimensional fused representation, where $d_{fused}=d_{PLM}$. The final output of the fusion network also undergoes Layer Normalization.
Finally, the fused representation $\mathbf{h}_{fused,i}$ is fed into a classification head, another MLP (Linear -> ReLU -> Dropout -> Linear), to predict the node label probabilities:
$\hat{\mathbf{y}}_{i}=\text{Softmax}(\text{MLP}_{classifier}(\mathbf{h}_{fused,i}))$ (9)
where $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{|\mathcal{Y}|}$ contains the predicted probabilities for each class in $\mathcal{Y}$.
5 GMLM Training Procedure
The training of GMLM proceeds in two distinct stages: (1) GNN Contrastive Pretraining, focused on learning robust structural representations and initializing the GNN-specific mask token, followed by (2) End-to-End Fine-tuning, where the entire model is trained for the downstream node classification task. All training stages and experiments for all 3 datasets were conducted on a single Tesla V100 (32GB) and took about $\sim$30 minutes with a maximum VRAM usage of 3GB. We also provide the impact on embeddings across different stages of training in  Appendix D.
5.1 Stage 1: GNN Contrastive Pretraining
The primary goal of this stage is to pre-train the GNN branch (Section 4.2) and the learnable GNN-specific mask token embedding $\mathbf{e}_{mask}^{gnn}$ (Section 4.1) to capture meaningful structural information from the graph. This is achieved through a contrastive learning objective. During this stage the weights of the PLM are frozen, but become trainable during Stage 2 as described in Section 5.2.
5.1.1 Data Augmentation via Soft Masking
For each training iteration in this stage, we generate two augmented "views" of the input node features. Two distinct perturbation masks, $\mathbf{m}_{pert,1}$ and $\mathbf{m}_{pert,2}$, are created by randomly selecting a proportion $p_{gnn\_mask}$ (randomly sampled from a uniform distribution $\mathcal{U}(0.2,0.5)$ in each iteration) of all nodes in the graph $\mathcal{V}$ to be perturbed. This selection can be uniform random or based on heuristics like node degree (the code uses degree-based sampling if degrees are non-zero, otherwise random). Using these two masks and the soft masking mechanism (Equation 4) with an interpolation weight $\beta$ ($\beta=0.7$), we obtain two sets of perturbed node features, $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$.
5.1.2 Contrastive Objective
Both $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$ are fed through the same shared-weight GNN encoder (specifically, we produce two sets of node embeddings, $\mathbf{H}_{G,1}$ and $\mathbf{H}_{G,2}\in\mathbb{R}^{N\times d_{G}}$. We then apply the NT-Xent contrastive loss (Equation 3) with a temperature $\tau_{pretrain}$ ($\tau_{pretrain}=0.5$). The pretraining loss $\mathcal{L}_{pretrain}$ is:
$\mathcal{L}_{pretrain}=\mathcal{L}_{NT-Xent}(\mathbf{H}_{G,1},\mathbf{H}_{G,2},\tau_{pretrain})$ (10)
We conduct contrastive pre-training for 30 epochs. During this pretraining phase, only the parameters of the GNN branch and $\mathbf{e}_{mask}^{gnn}$ are updated using an AdamW optimizer. Specifically, the GNN branch parameters (RGCN layers, batch normalization, residual projections) use a learning rate $lr_{gnn\_branch\_pretrain}$ ($1\times 10^{-3}$), while the GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$ uses a learning rate $lr_{token\_pretrain}$ ($1\times 10^{-4}$). Both use a weight decay of $0.01$. A Cosine Annealing learning rate scheduler is also employed.
5.2 Stage 2: End-to-End Fine-tuning for Node Classification
After GNN pretraining, the entire GMLM architecture, with the GNN branch initialized from Stage 1, is fine-tuned for node classification. We utilize the raw textual information associated with heterophilic graphs as input to the PLM. Such information is only available for three datasets (Cornell, Texas, and Wisconsin) which we obtain from Zenodo  [^2] https://zenodo.org/records/7555368.
5.2.1 Active Node Selection for PLM and Loss Computation
For each training iteration, an active node mask $\mathbf{m}_{active}$ is generated. This mask first considers nodes in the current training split ($\mathcal{V}_{train}$). From this set, a random proportion $p_{active\_nodes}$ (randomly sampled from a uniform distribution $\mathcal{U}(0.3,0.8)$ in each iteration) is selected based on node degree (or randomly if all degrees are zero). Only texts $\{t_{i}:m_{active,i}=1\}$ are processed by the PLM, and the classification loss is computed only over these active nodes.
5.2.2 GNN Input Perturbation during Fine-tuning
The soft masking mechanism (Equation 4) with the same interpolation weight $\beta$ ($\beta=0.7$) is also applied to the GNN input features during fine-tuning. The perturbation mask used is $\mathbf{m}_{active}$, meaning nodes whose texts are processed by the PLM also have their GNN input features (partially) masked by $\mathbf{e}_{mask}^{gnn}$ (which is also being fine-tuned).
5.2.3 Supervised Classification Loss
The model processes inputs to produce predicted label probabilities $\hat{\mathbf{y}}_{i}$. The supervised classification loss $\mathcal{L}_{classify}$ is the cross-entropy loss computed over active nodes $\mathcal{V}_{active}$. Label smoothing (with a factor of 0.1) is applied to the true labels $y_{i,c}$.
5.2.4 Optimization and Hyperparameters
The total loss $\mathcal{L}_{classify}$ is backpropagated. All trainable parameters are updated using the AdamW optimizer [23]. We employ differential learning rates:
  • GNN components (RGCN layers, residual projections, batch normalization): $lr_{gnn}$ ($1\times 10^{-3}$).
  • PLM (GTE-small & DistilBERT) parameters: $lr_{plm}$ ($1\times 10^{-5}$).
  • Other components (fusion network, classifier, GNN projection, GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$): $lr_{other}$ ($1\times 10^{-4}$).
A weight decay of $0.01$ is applied to all parameter groups. The model is trained for a maximum of 1000 epochs. We employ an early stopping strategy with a patience of 15 epochs: training halts if the F1 score on the validation set does not improve. The model state corresponding to the best validation F1 score is saved and used for final evaluation on the test set. The specific values for hyperparameters were set as follows (selected through a grid search the details of which are provided in  Appendix C):
  • $p_{gnn\_mask}$ (GNN Perturbation Ratio for Pretraining): During GNN contrastive pretraining, for each of the two generated views, the proportion of nodes whose input features are soft-masked is randomly sampled from a uniform distribution $\mathcal{U}(0.2,0.5)$ in each epoch.
  • $p_{active\_nodes}$ (Active Node Ratio for Fine-tuning): In each epoch of the end-to-end fine-tuning stage, the proportion of training nodes selected to be "active" (for PLM processing, GNN input perturbation, and loss computation) is randomly sampled from a uniform distribution $\mathcal{U}(0.3,0.8)$.
  • $\beta$ (Soft Masking Interpolation Weight): The interpolation weight for the soft masking mechanism (Equation 4) is set to $\beta=0.7$. This value is used during both the GNN pretraining stage and for GNN input feature perturbation in the fine-tuning stage.
  • $\tau_{pretrain}$ (Pretraining Contrastive Temperature): The temperature hyperparameter for the NT-Xent contrastive loss (Equation 3) during the GNN pretraining stage is set to $\tau_{pretrain}=0.5$.
Cornell Wisconsin Texas
MLP-2 91.30 $\pm$ 0.70 93.87 $\pm$ 3.33 92.26 $\pm$ 0.71
GAT[4] 76.00 $\pm$ 1.01 71.01 $\pm$ 4.66 78.87 $\pm$ 0.86
APPNP[24] 91.80 $\pm$ 0.63 92.00 $\pm$ 3.59 91.18 $\pm$ 0.70
GPRGNN[25] 91.36 $\pm$ 0.70 93.75 $\pm$ 2.37 92.92 $\pm$ 0.61
H2GCN[26] 86.23 $\pm$ 4.71 87.5 $\pm$ 1.77 85.90 $\pm$ 3.53
MixHop[27] 60.33 $\pm$ 28.53 77.25 $\pm$ 7.80 76.39 $\pm$ 7.66
GCN+JK 66.56 $\pm$ 13.82 62.50 $\pm$ 15.75 80.66 $\pm$ 1.91
GAT+JK 74.43 $\pm$ 10.24 69.50 $\pm$ 3.12 75.41 $\pm$ 7.18
FAGCN [28] 88.03 $\pm$ 5.6 89.75 $\pm$ 6.37 88.85 $\pm$ 4.39
BernNet[29] 92.13 $\pm$ 1.64 NA 93.12 $\pm$ 0.65
GraphSAGE[30] 71.41 $\pm$ 1.24 64.85 $\pm$ 5.14 79.03 $\pm$ 1.20
Geom-GCN*[31] 60.81 64.12 67.57
SGC-1 [32] 70.98 $\pm$ 8.39 70.38 $\pm$ 2.85 83.28 $\pm$ 5.43
SGC-2 [32] 72.62 $\pm$ 9.92 74.75 $\pm$ 2.89 81.31 $\pm$ 3.3
GCNII [33] 89.18 $\pm$ 3.96 83.25 $\pm$ 2.69 82.46 $\pm$ 4.58
GCNII+ [33] 90.49 $\pm$ 4.45 89.12 $\pm$ 3.06 88.52 $\pm$ 3.02
GCN [34] 82.46 $\pm$ 3.11 75.5 $\pm$ 2.92 83.11 $\pm$ 3.2
Snowball-2 82.62 $\pm$ 2.34 74.88 $\pm$ 3.42 83.11 $\pm$ 3.2
Snowball-3 82.95 $\pm$ 2.1 69.5 $\pm$ 5.01 83.11 $\pm$ 3.2
GMLM(GTE-Small) 96.74 $\pm$ 0.72 91.54 $\pm$ 2.23 94.83 $\pm$ 1.31
GMLM(DistilBERT) 96.87 $\pm$ 0.57 92.32 $\pm$ 2.21 95.16 $\pm$ 1.27
Table 1: Mean Test Accuracy(%) $\pm$ standard deviation on three heterophilic node classification tasks across 10 runs. Highest score is denoted by bold and second highest by underline. $\ast$ Denotes the best performing variant of multiple variants reported by the authors.
6 Results
We evaluated GMLM on three widely recognized heterophilic WebKB benchmark datasets [31]: Cornell, Wisconsin, and Texas. These datasets, involving the classification of university web pages (e.g., student, faculty), are particularly suitable for assessing models on heterophilic graphs (see Appendix A, Table 3 for details).
Our evaluation protocol involved 10 random 60%/20%/20% train/validation/test splits per dataset, reporting mean test accuracy and standard deviation for reproducibility. Baseline models were configured with their reported optimal hyperparameters. GMLM was evaluated with two PLM backbones (GTE-Small and DistilBERT), following the two-stage training procedure from Section 5. Table 1 compares the performance of GMLM variants against a suite of GNN baselines, highlighting the top score in bold and the second highest with an underline.
6.1 Performance on Cornell
On the Cornell dataset, GMLM demonstrates exceptional performance. Specifically, GMLM(DistilBERT) achieves the highest mean test accuracy of 96.87% $\pm$ 0.57, while GMLM(GTE-Small) secures the second-highest accuracy at 96.74% $\pm$ 0.72. Both variants significantly outperform all other baselines. The next best performing baseline is BernNet, which attains 92.13% $\pm$ 1.64, followed by GPRGNN at 91.36% $\pm$ 0.70. Compared to BernNet, GMLM(DistilBERT) shows an improvement of over 4.7%, highlighting GMLM’s superior ability to integrate textual and structural information on this heterophilic graph. The low standard deviations for both GMLM variants indicate stable performance across different random splits.
6.2 Performance on Wisconsin
On the Wisconsin dataset, the highest accuracy is achieved by MLP-2 at 93.87% $\pm$ 3.33, with GPRGNN closely following at 93.75% $\pm$ 2.37. Our GMLM(DistilBERT) variant scores 92.32% $\pm$ 2.21, ranking as the third-best performing model overall, while GMLM(GTE-Small) achieves 91.54% $\pm$ 2.23. Although GMLM does not claim the top position on this dataset, both variants outperform a majority of the other heterophilic GNN baselines, such as APPNP (92.00% $\pm$ 3.59) and H2GCN (87.5% $\pm$ 1.77). The performance of GMLM(DistilBERT) is competitive, being only 1.55% behind the top-performing MLP-2, showcasing its robustness even when simpler models perform strongly.
6.3 Performance on Texas
On the Texas dataset, GMLM again demonstrates state-of-the-art performance. GMLM(DistilBERT) achieves the highest mean test accuracy of 95.16% $\pm$ 1.27, and GMLM(GTE-Small) follows as the second-highest with 94.83% $\pm$ 1.31. Both GMLM variants substantially outperform all other baselines. The next best baseline is BernNet, which attains 93.12% $\pm$ 0.65, followed by GPRGNN at 92.92% $\pm$ 0.61. GMLM(DistilBERT) surpasses BernNet by over 2%. These results underscore GMLM’s effectiveness in handling heterophilic graph structures on the Texas dataset, with consistent performance indicated by their low standard deviations.
7 Timing Comparisons
Table 2 highlights the computational efficiency of our GMLM framework compared to a naive strategy that encodes all node texts at every training iteration with GTE-Small. Across all datasets, GMLM achieves consistent training time reductions of approximately 33% to 40%, attributable to its dynamic active node selection strategy. Additionally, GMLM exhibits marginally lower peak GPU memory usage, with savings of around 6% to 7%, validating its scalability advantage without compromising classification performance.
Dataset Method Pretrain (s) Fine-tune (s) Peak Mem (GB)
Cornell GMLM 1.436 3.081 1.25
Naïve 2.738 3.997 1.33
Texas GMLM 1.534 3.252 1.26
Naïve 2.883 4.012 1.34
Wisconsin GMLM 1.612 3.447 1.26
Naïve 2.994 4.102 1.35
Table 2: Comparison of GMLM vs. a naïve PLM+GNN integration on runtime (s/epoch) and GPU memory usage (GB).
8 Conclusion
This paper presents GMLM (Graph Masked Language Model), an innovative architecture aimed at seamlessly and effectively merging graph structural details with comprehensive textual meanings for node representation learning, especially in heterophilic graphs. GMLM tackles major obstacles in this area through three primary contributions: (i) a dynamic active node selection strategy that lessens the computational load of handling large text datasets by selectively activating the PLM; (ii) a GNN-specific contrastive pretraining phase employing soft masking with a trainable graph [MASK] token, which prepares the GNN for strong structural feature extraction; and (iii) a specialized fusion network that adaptively merges the generated GNN and PLM embeddings. Our comprehensive tests on popular heterophilic benchmark datasets (Cornell, Texas, and Wisconsin) confirm the effectiveness of GMLM. The findings indicate that GMLM, especially when employing PLMs such as DistilBERT and GTE-Small, attains either state-of-the-art or significantly competitive performance against a wide range of established GNN baselines. For example, GMLM(DistilBERT) achieved peak accuracies of 96.87% on Cornell and 95.16% on Texas, greatly surpassing previous approaches. Even with the Wisconsin dataset, where a basic MLP variant excelled, GMLM stayed competitive, highlighting its resilience. The reliable outcomes over several attempts and the small standard deviations emphasize the robustness of our method.
9 Limitations
The GMLM architecture, by design, utilizes textual information associated with graph nodes. This introduces a dependency on the nature of the input dataset, which constitutes the primary limitation of the current study. The effectiveness of GMLM is contingent upon the availability of raw textual data for the nodes. This work targets graphs with node-level raw textual content, which is essential for leveraging the PLM component in GMLM. Our choice of the WebKB datasets (Cornell, Texas, Wisconsin) is deliberate, in many widely used citation or product graphs (e.g., Cora, Citeseer, PubMed), node features are limited to bag-of-words vectors or preprocessed embeddings, and the original raw text is unavailable. Without node text, GMLM reduces to a pure GNN.
Appendix A Dataset Statistics
Table 3 provides an overview of the key statistics for the datasets used in our experiments.
Name Nodes Edges Features Classes
Cornell 183 298 1,703 5
Texas 183 325 1,703 5
Wisconsin 251 515 1,703 5
Table 3: Statistics of the heterophilic benchmark datasets used.
Appendix B Model Sizes
Table 4 details the approximate number of trainable parameters for the main components of our GMLM architecture. The RGCN branch size is calculated based on the configuration used for the WebKB datasets (Cornell, Texas, Wisconsin) which have an input feature dimension of 1,703 and a GNN hidden channel size of 512.
Component Parameters (Approx.)
GTE-Small (PLM) 33M
DistilBERT (PLM) 66M
RGCN Branch (GNN) $\sim$18.88M
Calculated for GNN input features $d_{x}=1703$, hidden channels $d_{h}=512$, 3 RGCN layers,
   and ‘num_relations=4‘. Includes GNN layers, batch norms, residual projections, and
   the GNN-specific mask token.
Table 4: Approximate model sizes (number of trainable parameters).
Appendix C Hyperparameter Ablation
This section provides further details on the hyperparameter tuning process for our GMLM architecture and an ablation analysis of key components. To identify the optimal configuration, we performed a grid search over several critical hyperparameters. The specific parameter ranges and values explored during this search are detailed in Table 5.
Figure 3 visualizes the performance (mean test-set accuracy across 10 runs using DistilBERT as the PLM backbone) of various hyperparameter configurations using a parallel coordinates plot. Each line in the plot represents a unique combination of hyperparameters, allowing for an intuitive understanding of how different settings influence the final accuracy on the benchmark datasets (Cornell, Wisconsin, and Texas). The axes correspond to the hyperparameters: soft masking interpolation weight ($\beta$), the range for active node selection ratio ($p_{active\_nodes}$), the range for GNN input perturbation ratio ($p_{gnn\_mask}$), and the contrastive pretraining temperature ($\tau_{pretrain}$). The plot highlights the configuration that yielded the best overall performance, as indicated by the thicker orange line and specified in the legend as (0.7, 0.55, 0.35, 0.7), where 0.55 and 0.35 represent the midpoints of the selected ranges for $p_{active\_nodes}$ (0.3, 0.8) and $p_{gnn\_mask}$ (0.2, 0.5) respectively. This visualization aids in understanding the sensitivity of GMLM to these hyperparameters and validates the choices made for the final reported results.
Appendix D Impact of Training Stages
To qualitatively assess the impact of different training stages on the learned node representations, Figure 2 visualizes the evolution of fused node embeddings for the Cornell dataset, using DistilBERT as the PLM backbone. The figure presents t-SNE (left column) and PCA (right column) projections of these embeddings at three key stages: before any training, after the GNN contrastive pretraining stage, and after complete end-to-end fine-tuning. Node colors in the plots correspond to their respective ground-truth classes.
Figure 2: Evolution of node embeddings on the Cornell dataset (DistilBERT PLM) visualized using t-SNE (left column) and PCA (right column) at different training stages: (Top Row) Before Training, (Middle Row) After GNN Contrastive Pretraining, and (Bottom Row) After Complete End-to-End Fine-tuning. Colors represent distinct node classes (0-4). The variance captured by the first two principal components is noted for PCA plots (e.g., "Var: 0.34").
Before Training (Top Row): Initially, before any training, the t-SNE plot shows that node embeddings from different classes are largely intermingled and lack clear separation. Similarly, the PCA visualization (Var: 0.34) indicates that the initial features, likely dominated by the raw input features and randomly initialized GNN/PLM components not yet fine-tuned, capture limited class-specific variance in the principal components. There is no discernible clustering according to node labels.
After Pretraining (Middle Row): Following the GNN contrastive pretraining stage, we observe some emergent structuring in the embedding space. Both t-SNE and PCA plots show early signs of clustering, with some same-colored points beginning to group together. This suggests that the contrastive pretraining has started to instill structurally relevant information into the GNN embeddings, leading to initial differentiation. However, the class separation is still not well-defined, and considerable overlap remains. The PCA plot (Var: 0.40) shows a slight increase in the explained variance by the top two components, aligning with this initial structural learning.
After Complete Training (Bottom Row): After the end-to-end fine-tuning phase, where both the GNN and the DistilBERT PLM are trained jointly for the node classification task, there is a marked improvement in class separability. The t-SNE visualization clearly shows more distinct and compact clusters for the different node classes. The PCA plot (Var: 0.61) now captures a substantially higher proportion of the variance in its first two principal components, reflecting the enhanced class structure in the embedding space. This indicates that the GMLM model, through its two-stage training and fusion mechanism, has learned to produce highly discriminative embeddings that are well-suited for the downstream node classification task.
Figure 3: Mean test-set accuracy across 10 runs for different hyperparameter configurations using DisitilBERT as the PLM. The averaged values such as 0.55 and 0.35 refer to (0.3,0.8) and (0.2,0.5) respectively.
Soft Mask $\boldsymbol{\beta}$ Active Node Ratio GNN Perturb. Ratio Contrastive Temp. $\boldsymbol{\tau}$
0.5 (0.2, 0.6) (0.1, 0.3) 0.3
0.7 (0.3, 0.8) (0.2, 0.5) 0.5
0.9 (0.4, 0.9) (0.3, 0.6) 0.7
Table 5: Grid search parameters explored for GMLM hyperparameter tuning. Ranges for ratio-based parameters are specified as (min, max) for uniform sampling during training.
Appendix E Method Comparison with other integrations of PLMs and GNNs
To further contextualize our proposed GMLM architecture, this section provides a detailed comparative analysis against prominent existing methods in graph-based neural network pre-training and integration with language models. Table 6 (from the main paper, re-iterated here for completeness or referenced if preferred) summarizes the key characteristics, objectives, and methodological differences between GMLM and influential works such as GPT-GNN [18], GraSAME [13], and the Graph Language Model (GLM) framework by Plenz and Frank [12]. This comparison highlights the unique aspects of GMLM, including its specific GNN contrastive pre-training strategy, the dynamic active node selection for efficient PLM engagement, and the explicit fusion mechanism for combining structural and textual representations, particularly for the task of heterophilic node classification.
Feature GPT-GNN GraSAME GLM GMLM (Ours)
Primary Goal Generative pre-training of GNNs for transfer learning. Inject token-level graph structure into PLMs for graph-to-text generation. Enable LMs to natively process graphs & text as Graph Transformers. Integrate GNN (structure) & PLM (rich node text) for heterophilic node classification.
Pre-training Focus GNN pre-training (generative: attribute & edge reconstruction). Relies on pre-trained PLM. GNN component trained during fine-tuning (PLM frozen). Initializes from pre-trained LM; explores LM’s adaptability to graph structure. GNN contrastive pre-training (soft masking, learnable [MASK] token). PLM is pre-trained & then fine-tuned.
GNN Component Any standard GNN (e.g., HGT, GCN). GNN (SAGE, GAT, RGCN) operates on token-level graph derived from linearized input. LM architecture itself is modified to become a Graph Transformer (GT). Dedicated RGCN branch operating on the input graph structure.
PLM Component Not a primary focus for rich node text; can use simpler models (e.g., LSTM) for text attributes. Standard PLM (e.g., T5 encoder), parameters are frozen during fine-tuning. LM (e.g., T5) architecture is modified to become a GT; parameters initialized from PLM. Standard PLM (e.g., GTE-Small, DistilBERT) for node text; parameters are fine-tuned.
Integration/ Fusion Method GNN learns attribute-structure dependency implicitly via the generative pre-training task. GNN output directly guides the Q vector in PLM’s self-attention mechanism. Transformer self-attention & positional encoding are adapted for graph processing within the LM architecture. GNN & PLM embeddings are explicitly concatenated & fused via a dedicated MLP.
Text/Attribute Handling Focuses on "node attributes" (can be vectors or generated text). Graph is linearized to text; GNN processes a token-level graph derived from this. Triplet components (textual) are processed by the GT-adapted LM. PLM processes raw textual descriptions of selected nodes. GNN uses original node features (potentially perturbed).
Scalability Strategy Subgraph sampling and an adaptive queue for large-scale graph pre-training. Reduces trainable parameters significantly by freezing the PLM. Not a primary focus of the presented work. Dynamic active node selection for PLM processing and loss computation.
Table 6: Comparison of GMLM with related language model integrations with graph neural networks. GMLM distinguishes itself through its specific GNN contrastive pre-training, dynamic active node selection for PLM engagement, and explicit fusion of separately processed GNN and PLM embeddings for node classification.
Appendix F LLM Usage
We used the Gemini2.5 Pro models to help refine our grammar and paper presentation.