### v4

# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

Abstract
Integrating powerful but computationally expensive Pre-trained Language Models (PLMs) with Graph Neural Networks (GNNs) is a key challenge, especially on text-rich heterophilic graphs. We propose the Graph Masked Language Model (GMLM), a framework designed for the efficient and effective fusion of graph structure and text semantics. GMLM employs a two-stage process: first, a contrastive pre-training stage with a novel soft masking technique builds a robust multi-scale GNN; second, an end-to-end fine-tuning stage uses a dynamic active node selection strategy for scalability and a bi-directional cross-attention module for deep fusion. Experiments on five heterophilic benchmarks show GMLM achieves state-of-the-art results on four, significantly outperforming prior GNN and large LLM-based methods. For instance, it improves accuracy on the Texas dataset by over 8% and on Wisconsin by nearly 5%. Our work demonstrates that a sophisticated, deeply-integrated architecture can be more effective and efficient than larger, general-purpose models for text-rich graph representation learning.
GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification
Aarush Sinha Independent Researcher aarush.sinha@gmail.com
1 Introduction
Graph Neural Networks (GNNs) Scarselli et al. (2009); Zhou et al. (2021) have demonstrated remarkable success in node classification, link prediction, and graph-level tasks. However, many such architectures focus primarily on structural information and local neighborhood features, potentially overlooking rich semantic content embedded within node attributes, especially textual attributes. While recent advances in natural language processing, particularly Pre-trained Language Models (PLMs) like BERT Devlin et al. (2019), have revolutionized our ability to capture contextual information from text, their effective and scalable integration with graph learning remains an active area of research. The challenge is particularly pronounced in heterophilic graphs, where connected nodes may exhibit dissimilar features and labels, requiring a nuanced understanding of both structure and semantics.
While architectures like Graph Attention Networks (GATs) Veličković et al. (2018) offer learnable attention mechanisms for neighborhood aggregation, and Relational Graph Convolutional Networks (RGCNs) Schlichtkrull et al. (2017) can handle typed edges, they do not inherently perform deep semantic reasoning on node-associated texts without explicit integration with PLMs. Concurrently, masked modeling techniques, central to the success of PLMs, have proven highly effective for learning contextual representations. This prompts the question: How can we develop a deeply integrated framework that combines the strengths of GNNs and PLM-style masked modeling to enhance representation learning on text-rich graphs, while addressing computational scalability?
Traditional node feature masking in GNNs Mishra et al. (2021) often involves random feature perturbation or complete removal, which can disrupt learning or fail to provide sufficiently informative training signals. Furthermore, applying PLMs to all node texts in large graphs is often computationally prohibitive.
In this paper, we present a novel framework that deeply integrates GNNs and PLMs for improved node representation learning on text-rich graphs, with a focus on heterophilic settings and computational efficiency. Our approach introduces three key innovations:
  1. (1)
    A dynamic active node selection strategy that stochastically identifies a subset of nodes during each training iteration. By processing only the texts of these active nodes with the PLM, we drastically reduce the computational overhead and memory footprint, making the approach scalable.
  2. (2)
    A GNN-specific contrastive pre-training stage that pre-trains a deep, multi-scale GNN (an RGCN in our implementation) to learn robust structural representations. This stage employs a novel soft masking mechanism that interpolates original node features with a learnable, graph-specific [MASK] token embedding, providing a more stable and informative training signal than hard masking.
  3. (3)
    A bi-directional cross-attention fusion module that moves beyond simple concatenation. It allows the multi-scale graph embeddings and the deep semantic PLM embeddings to mutually inform and refine one another, creating a rich, contextually-aware joint representation before the final classification.
This combination of selective PLM processing, robust GNN pre-training, and sophisticated fusion allows our model to learn both structural and semantic patterns effectively, leading to more powerful and contextually aware node representations for challenging node classification tasks.
Our main contributions in this work can be summarized as follows:
  1. 1.
    We propose a novel and scalable framework that deeply integrates a multi-scale GNN with a PLM. The framework’s efficiency is achieved through a dynamic active node selection mechanism for on-demand text processing, making it tractable for large graphs.
  2. 2.
    We introduce a GNN-specific contrastive pre-training stage that utilizes a soft feature masking technique with a learnable graph [MASK] token. This stage is designed to effectively train a deep, multi-scale GNN architecture, allowing it to learn robust and hierarchical structural representations before end-to-end fine-tuning (Section 5.1).
  3. 3.
    We design a sophisticated bi-directional cross-attention fusion module that moves beyond simple concatenation. This mechanism allows the graph and text representations to mutually inform and refine one another, creating a richer, more contextually-aware joint representation for improved node classification, especially on challenging heterophilic graphs (Section 5.2).
2 Related Works
Graph Neural Networks (GNNs) Zhou et al. (2021) provide a powerful framework for learning from graph-structured data. Building upon early spectral and spatial convolutional methods, Graph Attention Networks (GATs) Veličković et al. (2018) integrated self-attention Vaswani et al. (2017). This allows adaptive weighting of neighbors during message passing, enhancing model expressivity and suitability for irregular graph structures, as detailed in recent surveys.
Efforts to adapt the Transformer architecture, initially designed for sequences, to graph data aim to capture long-range dependencies via global self-attention, potentially overcoming limitations of traditional message passing Ying et al. (2021), Yun et al. (2020). These Graph Transformers, reviewed systematically Shehzad et al. (2024), can incorporate positional information, handle diverse graph scales, achieve competitive results, and utilize language model pretraining benefits on text-rich graphs.
Masked Pretraining Strategies, popularized by language models Devlin et al. (2019) that learn contextual representations via masked token prediction, have been adapted for graphs. The Masked Graph Autoencoder (MaskGAE) Li et al. (2023a) framework, for instance, masks graph edges and trains the model to reconstruct them, thereby learning robust representations for downstream tasks like link prediction and node classification.
Integration of Language Models and Graph Structures seeks to combine textual and structural information. Key directions include: fusing pretrained language models (LMs) with GNNs for joint reasoning over text and structure Plenz and Frank (2024); injecting structural information directly into LMs using graph-guided attention mechanisms, potentially removing the need for separate GNN modules Yuan and Färber (2024); interpreting Transformers as GNNs operating on complete graphs, allowing explicit modeling of edge information Henderson et al. (2023); and jointly training large LMs and GNNs to utilize both contextual understanding and structural learning capabilities Ioannidis et al. (2022).
3 Preliminaries
In this section, we introduce the fundamental concepts and notations used throughout this paper, including graph representation, Graph Neural Networks (specifically RGCNs), Pre-trained Language Models, and the concept of heterophily in graphs.
3.1 Graph Representation and Notation
A graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of $N=|\mathcal{V}|$ nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$ is the set of $M=|\mathcal{E}|$ edges. Each edge $(v_{i},r,v_{j})\in\mathcal{E}$ connects node $v_{i}$ to node $v_{j}$ under a relation type $r\in\mathcal{R}$, where $\mathcal{R}$ is the set of possible relation types. For graphs without explicit multiple relation types, we can consider a single relation type, effectively making it $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Each node $v_{i}\in\mathcal{V}$ is associated with an initial feature vector $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$, where $d_{x}$ is the input feature dimensionality. In the context of this work, nodes are also associated with textual descriptions $T=\{t_{1},t_{2},\dots,t_{N}\}$, where $t_{i}$ is the raw text for node $v_{i}$. The goal of node classification is to predict a label $y_{i}\in\mathcal{Y}$ for each node (or a subset of nodes), where $\mathcal{Y}$ is the set of possible labels.
3.2 Graph Neural Networks (GNNs)
Graph Neural Networks (GNNs) are a class of neural networks designed to operate directly on graph-structured data. They typically learn node representations by aggregating information from their local neighborhoods. A generic GNN layer can be expressed as:
$\mathbf{h}_{i}^{(l+1)}=\text{UPDATE}^{(l)}\left(\mathbf{h}_{i}^{(l)},\text{AGGREGATE}^{(l)}\left(\{(\mathbf{h}_{j}^{(l)},\mathbf{e}_{ij}):v_{j}\in\mathcal{N}(v_{i})\}\right)\right)$
where $\mathbf{h}_{i}^{(l)}$ is the representation of node $v_{i}$ at layer $l$, $\mathcal{N}(v_{i})$ is the set of neighbors of $v_{i}$, $\mathbf{e}_{ij}$ is the feature of the edge between $v_{i}$ and $v_{j}$ (if available), and AGGREGATE and UPDATE are differentiable functions ( sum, mean, max, or more complex neural networks).
3.2.1 Relational Graph Convolutional Networks (RGCNs)
Relational Graph Convolutional Networks (RGCNs) Schlichtkrull et al. (2017) extend GCNs to handle multi-relational graphs (heterogeneous graphs with different edge types). The message passing for a node $v_{i}$ in an RGCN layer is typically defined as:
$\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{v_{j}\in\mathcal{N}_{r}(v_{i})}\frac{1}{c_{i,r}}\mathbf{W}_{r}^{(l)}\mathbf{h}_{j}^{(l)}+\mathbf{W}_{0}^{(l)}\mathbf{h}_{i}^{(l)}\right)$ (2)
where $\mathcal{N}_{r}(v_{i})$ is the set of neighbors of node $v_{i}$ under relation $r\in\mathcal{R}$, $\mathbf{W}_{r}^{(l)}$ is a relation-specific learnable weight matrix for relation $r$ at layer $l$, $\mathbf{W}_{0}^{(l)}$ is a weight matrix for self-loops, $c_{i,r}$ is a normalization constant ( $|\mathcal{N}_{r}(v_{i})|$), and $\sigma$ is an activation function ( ReLU). To handle a large number of relations, RGCNs often use basis decomposition or block-diagonal decomposition for $\mathbf{W}_{r}^{(l)}$.
3.3 Pre-trained Language Models (PLMs)
Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2019) and its variants ( DistilBERT Sanh et al. (2020)), have achieved state-of-the-art results on various natural language processing (NLP) tasks. These models are typically based on the Transformer architecture Vaswani et al. (2023) and are pre-trained on massive text corpora using self-supervised objectives like Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). Given an input text sequence ( a node’s textual description $t_{i}$), a PLM first tokenizes it into a sequence of sub-word tokens $[w_{1},w_{2},\dots,w_{S}]$. These tokens are then converted into input embeddings (sum of token, position, and segment embeddings) and fed into multiple Transformer layers. The output of the PLM is a sequence of contextualized embeddings for each token. For sentence-level or document-level tasks, often the embedding of a special [CLS] token ( $\mathbf{h}_{[CLS]}$) or an aggregation ( mean pooling) of token embeddings is used as the representation of the entire input text. In our work, we utilize a PLM (specifically DistilBERT) to obtain semantic embeddings $\mathbf{h}_{PLM}(t_{i})$ from node texts $t_{i}$.
3.4 Node Feature Masking in GNNs
Masking is a common technique in self-supervised learning. In the context of GNNs, node feature masking aims to train the model to reconstruct or predict properties of masked nodes or their features, thereby learning robust representations. A simple approach might involve replacing a fraction of node features $\mathbf{x}_{i}$ with a zero vector or a special [MASK] token embedding before feeding them into the GNN Hu et al. (2020); You et al. (2020). Our GNN contrastive pretraining phase builds upon this idea but introduces a soft masking mechanism with a learnable graph-specific [MASK] token.
3.5 Heterophily in Graphs
Graph homophily refers to the principle that nodes tend to connect to other nodes that are similar to themselves ( in terms of features or labels). Conversely, graph heterophily (or disassortativity) describes the scenario where connected nodes are often dissimilar Zhu et al. (2020a). Many real-world graphs exhibit heterophily, such as protein-protein interaction networks or certain social networks. Standard GNNs, which assume homophily by smoothing features over neighborhoods, may perform poorly on heterophilic graphs. Designing GNN architectures that can effectively handle heterophily is an ongoing research challenge, often requiring models to capture more complex relational patterns beyond simple neighborhood similarity. Our work aims to contribute to this by integrating rich textual semantics which can provide crucial context in such settings.
3.6 Contrastive Learning
Contrastive learning is a self-supervised learning paradigm that aims to learn representations by maximizing agreement between differently augmented "views" of the same data sample (positive pairs) while minimizing agreement between views of different samples (negative pairs). A common contrastive loss function is the NT-Xent loss Sohn (2016):
$\mathcal{L}_{NT\text{-}Xent}=-\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{j})/\tau)}{\sum_{k=1}^{2B}\mathbb{I}_{[k\neq i]}\exp(\text{sim}(\mathbf{z}_{i},\mathbf{z}_{k})/\tau)}$
where $(\mathbf{z}_{i},\mathbf{z}_{j})$ is a positive pair of augmented representations from the same original sample within a batch of size $B$, $\text{sim}(\cdot,\cdot)$ is a similarity function ( cosine similarity), $\tau$ is a temperature hyperparameter, and the sum in the denominator is over all $2B$ augmented samples in the batch, excluding $\mathbf{z}_{i}$ itself. Our GNN pretraining phase utilizes this principle.
Figure 1: Caption
4 Methodology
We introduce a novel architecture designed to effectively integrate deep graph structural information with rich textual semantics from nodes for robust node representation learning. The architecture is characterized by two main processing pathways for graph and text data, a multi-scale GNN with residual connections, a sophisticated bi-directional cross-attention fusion module, and specialized input perturbation mechanisms to enhance robustness and training stability.
4.1 GNN Input Feature Perturbation with Soft Masking
To enhance the robustness of the GNN and to provide a meaningful training signal, we employ a soft masking strategy for the input node features. This is applied during both a contrastive pre-training phase and the final end-to-end fine-tuning. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{x}}$ be the matrix of initial node features for $N$ nodes. Given a binary perturbation mask $\mathbf{m}_{pert}\in\{0,1\}^{N}$, the soft masking operation produces perturbed features $\mathbf{X}^{\prime}$ as follows:
$\mathbf{x}^{\prime}_{i}=\begin{cases}(1-\beta)\cdot\mathbf{x}_{i}+\beta\cdot\mathbf{e}_{mask}^{gnn}&\text{if }m_{pert,i}=1\\ \mathbf{x}_{i}&\text{if }m_{pert,i}=0\end{cases}$ (4)
where $\mathbf{x}_{i}$ is the original feature vector for node $v_{i}$, $\mathbf{e}_{mask}^{gnn}\in\mathbb{R}^{d_{x}}$ is a learnable GNN-specific mask token embedding, and $\beta\in[0,1]$ is a hyperparameter controlling the interpolation weight. This GNN-specific mask token $\mathbf{e}_{mask}^{gnn}$ is initialized using Xavier uniform initialization and learned alongside other model parameters.
4.2 Multi-Scale Relational GNN Branch
The GNN branch is designed as a deep, multi-scale architecture to capture structural information across different neighborhood depths. It utilizes a stack of four Relational Graph Convolutional Network (RGCN) layers. The input to the first RGCN layer is the perturbed node feature matrix $\mathbf{X}^{\prime}$ from Equation 4.
Each of the four GNN blocks consists of an RGCNConv layer, followed by GraphNorm, a GELU activation function, and Dropout. To facilitate the training of this deep architecture and combine features from different network depths, we incorporate residual connections with linear projections. The hidden representation $\mathbf{H}^{(l)}$ at layer $l$ is computed as:
$\mathbf{Z}^{(l)}$ $=\text{GELU}(\text{GraphNorm}(\text{RGCN}^{(l)}(\mathbf{H}^{(l-1)},\mathcal{E})))$ (5)
$\mathbf{H}^{(l)}$ $=\text{Dropout}(\mathbf{Z}^{(l)})+\text{Proj}^{(l-1)}(\mathbf{H}^{(l-2)})$ (6)
where $\mathbf{H}^{(0)}=\mathbf{X}^{\prime}$, $\mathcal{E}$ represents the graph edges and their types, and $\text{Proj}^{(l-1)}$ is a linear projection. This structure allows the model to maintain and propagate information from earlier layers. We employ gradient checkpointing [^1] Specifically, torch.utils.checkpoint.checkpoint is used for each GNN block to trade a small amount of re-computation for significant memory savings, enabling the use of a deeper GNN. on each GNN block to manage memory consumption during backpropagation.
Instead of using only the output of the final GNN layer, we adopt a multi-scale fusion strategy to create a more comprehensive graph representation. The embeddings from all four GNN layers, $\{\mathbf{H}^{(1)},\mathbf{H}^{(2)},\mathbf{H}^{(3)},\mathbf{H}^{(4)}\}$, are collected. These are then combined using a learnable weighted sum after being projected to the PLM’s hidden dimension, $d_{PLM}$:
$\mathbf{h}_{G,i}=\text{LayerNorm}\left(\sum_{l=1}^{4}w_{l}\cdot\text{Proj}^{(l)}(\mathbf{h}_{i}^{(l)})\right)$ (7)
where $\mathbf{h}_{i}^{(l)}$ is the embedding for node $v_{i}$ from layer $l$, $\text{Proj}^{(l)}$ is a linear projection to dimension $d_{PLM}$, and $w_{l}$ are learnable weights derived from a softmax-normalized parameter vector. The final graph embedding for node $v_{i}$ is $\mathbf{h}_{G,i}\in\mathbb{R}^{d_{PLM}}$.
4.3 Pre-trained Language Model (PLM) Branch
The PLM branch processes the textual descriptions associated with nodes to capture rich semantic information. To manage computational costs, especially on large graphs, we process texts in micro-batches. During each training iteration, an active node mask $\mathbf{m}_{active}\in\{0,1\}^{N}$ identifies a subset of nodes $\mathcal{V}_{active}$ whose texts will be processed.
For each node $v_{i}\in\mathcal{V}_{active}$, its text $t_{i}$ is tokenized and fed into a pre-trained language model, such as GTE-Base (110M)Li et al. (2023b) and Snowflake-Embed(305M)Yu et al. (2024). To derive a fixed-size representation from the PLM’s variable-length output, we perform mean pooling over the last hidden state, weighted by the attention mask to exclude padding tokens. The PLM embedding for an active node $v_{i}$ is:
$\mathbf{h}_{PLM,i}=\frac{\sum_{j=1}^{L}\mathbf{o}_{j}\cdot a_{j}}{\sum_{j=1}^{L}a_{j}}$ (8)
where $\mathbf{O}=\text{PLM-Encoder}(t_{i})$ is the last hidden state sequence of length $L$, $\mathbf{o}_{j}$ is the embedding of the $j$-th token, and $a_{j}$ is the corresponding value from the attention mask (1 for real tokens, 0 for padding). For inactive nodes $v_{k}\notin\mathcal{V}_{active}$, their PLM embeddings are set to zero vectors, resulting in a complete PLM embedding matrix $\mathbf{H}_{PLM}\in\mathbb{R}^{N\times d_{PLM}}$.
4.4 Bi-Directional Attention Fusion and Classification
To achieve a deep and meaningful integration of graph and text modalities, we move beyond simple concatenation and employ a bi-directional cross-attention mechanism. This allows the graph and text representations to mutually inform and refine one another. The full node embedding matrices, $\mathbf{H}_{G}$ (from Section 4.2) and $\mathbf{H}_{PLM}$ (from Section 4.3), serve as inputs.
The cross-attention mechanism computes two sets of attended features:
  1. 1.
    Graph-to-Text Attention: The PLM embeddings serve as queries to attend to the GNN embeddings, producing text-aware graph features, $\mathbf{H}_{G\to T}$.
  2. 2.
    Text-to-Graph Attention: The GNN embeddings serve as queries to attend to the PLM embeddings, producing structure-aware text features, $\mathbf{H}_{T\to G}$.
The standard scaled dot-product attention is used, where for a query matrix $\mathbf{Q}$, key matrix $\mathbf{K}$, and value matrix $\mathbf{V}$:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (9)
The two attended representations are then concatenated for each node:
$\mathbf{h}_{concat,i}=[\mathbf{h}_{G\to T,i}\,;\,\mathbf{h}_{T\to G,i}]\in\mathbb{R}^{2\cdot d_{PLM}}$ (10)
This concatenated vector is passed through a fusion network, which is a feed-forward network consisting of a linear layer, LayerNorm, a GELU activation, and Dropout, to produce the final integrated representation:
$\mathbf{h}_{fused,i}=\text{FusionNet}(\mathbf{h}_{concat,i})\in\mathbb{R}^{d_{fused}}$ (11)
Finally, this fused representation $\mathbf{h}_{fused,i}$ is fed into a classification head, another MLP (Linear -> GELU -> Dropout -> Linear), to predict the node label probabilities:
$\hat{\mathbf{y}}_{i}=\text{Softmax}(\text{MLP}_{classifier}(\mathbf{h}_{fused,i}))$ (12)
where $\hat{\mathbf{y}}_{i}\in\mathbb{R}^{|\mathcal{Y}|}$ contains the predicted probabilities for each class in the label set $\mathcal{Y}$.
5 Training Procedure
The model training proceeds in two distinct stages: (1) GNN Contrastive Pre-training, focused on learning robust structural representations and initializing the GNN-specific mask token, followed by (2) End-to-End Fine-tuning, where the entire model is trained for the downstream node classification task. This two-stage approach allows the GNN to develop a strong structural foundation before being integrated with the powerful but computationally intensive language model. We also provide an analysis of the impact on embeddings across different stages of training in LABEL:app:imp-train.
5.1 Stage 1: GNN Contrastive Pre-training
The primary goal of this stage is to pre-train the multi-scale GNN branch (Section 4.2) and the learnable GNN-specific mask token embedding $\mathbf{e}_{mask}^{gnn}$ (Section 4.1) to capture meaningful structural information. This is achieved through a contrastive learning objective. During this stage, the weights of the PLM encoder, fusion network, and final classifier are frozen.
5.1.1 Data Augmentation via Soft Masking
For each training iteration, we generate two augmented "views" of the graph. Two distinct perturbation masks, $\mathbf{m}_{pert,1}$ and $\mathbf{m}_{pert,2}$, are created by selecting a proportion $p_{gnn\_mask}$ (randomly sampled from $\mathcal{U}(0.2,0.4)$ each iteration) of all nodes in the graph. This selection is weighted by node degree to prioritize more connected nodes. Using these masks and the soft masking mechanism (Equation 4) with an interpolation weight $\beta=0.7$, we obtain two sets of perturbed node features, $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$.
5.1.2 Contrastive Objective
Both $\mathbf{X}^{\prime}_{1}$ and $\mathbf{X}^{\prime}_{2}$ are passed through the same shared-weight GNN encoder to produce two sets of multi-scale node embeddings, $\mathbf{H}_{G,1}$ and $\mathbf{H}_{G,2}$. We then apply the NT-Xent contrastive loss with a temperature $\tau_{pretrain}=0.1$. The pre-training loss is:
$\mathcal{L}_{pretrain}=\mathcal{L}_{NT-Xent}(\mathbf{H}_{G,1},\mathbf{H}_{G,2},\tau_{pretrain})$ (13)
This pre-training is conducted for 30 epochs. Only the parameters of the GNN branch (RGCN layers, graph normalization, residual projections, multi-scale fusion) and the mask token $\mathbf{e}_{mask}^{gnn}$ are updated. We use an AdamW optimizer with a learning rate of $1\times 10^{-4}$ for all trainable parameters in this stage and a weight decay of $0.05$. A CosineAnnealingWarmRestarts learning rate scheduler is employed to adjust the learning rate throughout pre-training.
5.2 Stage 2: End-to-End Fine-tuning for Node Classification
After GNN pre-training, the entire model, with the GNN branch initialized from Stage 1, is fine-tuned for node classification. The model is designed to leverage raw textual information associated with nodes, which is particularly valuable for navigating the complex relationships in heterophilic graphs.
5.2.1 Active Node Selection and GNN Perturbation
For each training iteration, an active node mask $\mathbf{m}_{active}$ is generated. This mask first considers nodes in the training split ($\mathcal{V}_{train}$). From this set, a random proportion $p_{active\_nodes}$ (sampled from $\mathcal{U}(0.3,0.8)$) is selected based on node degree. This active mask serves two purposes:
  1. 1.
    PLM Processing: Only the texts $\{t_{i}:m_{active,i}=1\}$ are processed by the PLM to manage computational load.
  2. 2.
    GNN Input Perturbation: The same mask $\mathbf{m}_{active}$ is used to apply soft masking (Equation 4) to the GNN’s input features, continuing to regularize the GNN during fine-tuning.
The final classification loss is also computed exclusively over these active nodes.
5.2.2 Supervised Classification Loss
The model processes the inputs to produce predicted label probabilities $\hat{\mathbf{y}}_{i}$ (Equation 12). The supervised classification loss $\mathcal{L}_{classify}$ is the cross-entropy loss with label smoothing (smoothing factor of 0.2) applied to the true labels.
5.2.3 Optimization and Hyperparameters
The total loss $\mathcal{L}_{classify}$ is backpropagated through the entire model. All trainable parameters are updated using the AdamW optimizer Loshchilov and Hutter (2019) with differential learning rates:
  • GNN components (RGCN layers, projections, etc.): $lr_{graph}=1\times 10^{-4}$.
  • PLM (GTE-Base) parameters: $lr_{bert}=1\times 10^{-5}$.
  • Other components (attention, fusion, classifier, mask token): $lr_{other}=1\times 10^{-4}$.
A weight decay of $0.05$ is applied to the GNN and "other" parameter groups, while the PLM uses a standard weight decay of $0.01$. The model is trained for a maximum of 500 epochs. A linear learning rate scheduler with a warm-up period corresponding to the first 10% of training steps is used. Additionally, gradient clipping with a maximum norm of 1.0 is applied to prevent exploding gradients. We employ an early stopping strategy with a patience of 30 epochs: training halts if the F1 score on the validation set does not improve. The model state with the best validation F1 score is used for the final evaluation.
The key hyperparameters are summarized below:
  • $p_{gnn\_mask}$ (Pre-training Perturbation Ratio): Sampled from $\mathcal{U}(0.2,0.4)$ in each pre-training epoch.
  • $p_{active\_nodes}$ (Fine-tuning Active Node Ratio): Sampled from $\mathcal{U}(0.3,0.8)$ in each fine-tuning epoch.
  • $\beta$ (Soft Masking Weight): Set to $0.7$ for both stages.
  • $\tau_{pretrain}$ (Pre-training Contrastive Temperature): Set to $0.1$.
Methods Cornell Texas Wisconsin Actor Amazon
Classic GNNs
GCN Kipf and Welling (2017) 52.86$\pm$1.8 43.64$\pm$3.3 41.40$\pm$1.8 66.70$\pm$1.3 39.33$\pm$1.0
GraphSAGEHamilton et al. (2017) 75.71$\pm$1.8 81.82$\pm$2.5 80.35$\pm$1.3 70.37$\pm$0.1 46.63$\pm$0.1
GAT 54.28$\pm$5.1 51.36$\pm$2.3 50.53$\pm$1.7 63.74$\pm$6.7 35.12$\pm$6.4
Heterophily-specific GNNs
H2GCNZhu et al. (2020b) 69.76$\pm$3.0 79.09$\pm$3.5 80.18$\pm$1.9 70.73$\pm$0.9 47.09$\pm$0.3
FAGCNBo et al. (2021) 76.43$\pm$3.1 84.55$\pm$4.8 83.16$\pm$1.4 75.58$\pm$0.5 49.83$\pm$0.6
JacobiConvWang and Zhang (2022) 73.57$\pm$4.3 81.80$\pm$4.1 76.31$\pm$11.3 73.81$\pm$0.3 49.43$\pm$0.5
GBK-GNNDu et al. (2022) 66.19$\pm$2.8 80.00$\pm$3.0 72.98$\pm$3.3 72.49$\pm$1.0 44.90$\pm$0.3
OGNNWang et al. (2022) 71.91$\pm$1.8 85.00$\pm$2.3 79.30$\pm$2.1 72.08$\pm$2.4 47.79$\pm$1.6
SEGSL 66.67$\pm$4.1 85.00$\pm$2.0 79.30$\pm$1.8 72.73$\pm$0.8 47.38$\pm$0.2
DisamGCLZhao et al. (2024) 50.48$\pm$2.0 65.00$\pm$1.2 57.89$\pm$0.0 67.78$\pm$0.3 43.90$\pm$0.4
LLM4HeG Wu et al. (2025) (fine-tuned LLM/SLMs and distilled SLMs)
Vicuna 7B 77.62$\pm$2.9 89.09$\pm$3.3 86.14$\pm$2.1 76.82$\pm$0.5 51.53$\pm$0.4
Bloom 560M 75.48$\pm$2.1 80.00$\pm$4.0 86.49$\pm$1.9 76.16$\pm$0.6 51.52$\pm$0.5
Bloom 1B 75.71$\pm$1.4 83.86$\pm$2.8 83.86$\pm$1.7 74.99$\pm$0.5 52.33$\pm$0.6
7B-to-560M 75.00$\pm$4.0 88.18$\pm$2.2 87.19$\pm$2.5 75.78$\pm$0.2 51.51$\pm$0.4
7B-to-1B 77.38$\pm$2.7 88.18$\pm$4.0 86.14$\pm$1.5 75.37$\pm$0.9 51.58$\pm$0.4
GMLM
GTE-Base 74.25$\pm$2.5 85.90$\pm$4.9 89.81$\pm$2.5 73.89$\pm$2.1 50.64$\pm$0.7
Snowflake-embed 79.16$\pm$4.8 97.18$\pm$2.9 92.04$\pm$1.6 76.81$\pm$0.9 54.63$\pm$0.4
Table 1: Accuracy for node classification of different methods. Bold denotes the best score while underline denotes the second best.
6 Results
We evaluate GMLM on five widely-used heterophilic benchmark datasets. Following standard protocols, we randomly split the nodes into train, validation, and test sets with a proportion of 48%/32%/20% for the Cornell, Texas, Wisconsin, and Actor datasets, and 50%/25%/25% for Amazon following Platonov et al. (2024). All reported results in Table 1 are the mean accuracy (%) and standard deviation averaged over 10 independent runs.
As detailed in the table, our GMLM framework achieves state-of-the-art performance on the majority of benchmarks. The GMLM-Snowflake-embed variant establishes new records on four of the five datasets: Cornell, Texas, Wisconsin, and Amazon. The performance gains are particularly striking on the Texas dataset, where our model achieves 97.18% accuracy, surpassing the strongest LLM-based baseline by over 8 absolute percentage points. Similarly, it improves the state-of-the-art on Wisconsin by nearly 5 points to 92.04%.
Crucially, GMLM consistently outperforms both specialized heterophily-GNNs and even the much larger Vicuna 7B model. This demonstrates that our architectural design, which deeply integrates a pre-trained GNN with a PLM via bi-directional cross-attention, is more effective and capital-efficient than simply applying a larger, general-purpose model. On the Actor dataset, our result is highly competitive, nearly matching the 7B model’s performance with a significantly smaller PLM backbone.
A comparison of our model variants reveals that performance scales with the representational power of the PLM, as Snowflake-embed (305M) consistently outperforms GTE-Base (110M). However, even the smaller GMLM-GTE-Base is highly competitive, outperforming the majority of all baselines and securing the second-best result on Wisconsin. This confirms the robustness and effectiveness of the core GMLM framework itself.
7 Conclusion
In this paper, we introduced the Graph Masked Language Model (GMLM), a novel framework that effectively and efficiently integrates deep graph structural learning with the rich semantic understanding of Pre-trained Language Models. By employing a two-stage training procedure featuring a contrastive GNN pre-training phase and an end-to-end fine-tuning stage with dynamic active node selection, our model addresses key challenges of scalability and robust representation learning. The use of a sophisticated bi-directional cross-attention module for fusion proved critical, enabling GMLM to achieve new state-of-the-art results on four out of five widely-used heterophilic benchmarks. Our model significantly outperforms both specialized GNNs and much larger LLM-based approaches, underscoring that a carefully designed, deeply integrated architecture can be more powerful and capital-efficient than brute-force scaling for text-rich graph representation learning.
8 Limitations and Future Work
Despite the strong performance of GMLM, we acknowledge the limitations that present opportunities for future research. The framework’s effectiveness is intrinsically tied to the availability and quality of textual descriptions for nodes. In scenarios where such texts are sparse, noisy, or entirely absent, the PLM branch offers diminished value. Future work could explore robust fallbacks or methods for generating informative pseudo-texts to enhance performance on text-poor graphs.
Appendix A Dataset Statistics
Table 2 provides an overview of the key statistics for the datasets used in our experiments from HeTGB Li et al. (2025).
Name Nodes Edges H($G$) Classes
Cornell 195 304 0.13 5
Texas 187 328 012 5
Wisconsin 265 530 0.20 5
Actor 4416 12172 0.56 5
Wisconsin 24492 93050 0.38 5
Table 2: Statistics of the heterophilic benchmark datasets used.
Appendix B LLM Usage
We used the Gemini2.5 Pro models to help refine our grammar and paper presentation.